# Evidence of the matrix effect on a compositionally graded oxide thin film

### v1

@@ I Introduction
This original growth technique offers an ideal ground for exploring the SIMS matrix effect. This effect designates variations of signal which are not related to the atomic concentration. The secondary ion intensity collected by the instrument indeed combines several factors including the concentration of the atom of interest in the bombarded material, its the sputtering yield and its ionization efficiency. The latter two depends on the chemical and morphological properties of the matrix. Therefore significant variations of the ionic intensity can occur from sample to sample independently to the atomic concentration which is further sought. Lots of efforts have been made to elucidate the effect of a matrix on the ionic yield of diluted traces. Although empirical trends have been identified very early [5], the vast variety of the material chemistry and of the analysis conditions have made the task tedious and little success to build a unified theory for all cases have been met so far [6]. A common approach to reach SIMS quantitative measurements consists of interpolating among chemically similar standards using semi-empirical models with limited validity domains [7]. This means to track ion yields in samples with similar compositions; it can be either series of alloys [8] or multi-implanted samples [9].
@@ II.2 SIMS conditions
The SIMS analyses were performed with a magnetic sector analyser (IMS7f-CAMECA). To avoid surface charging, the sample surface was coated with gold (50 nm) prior analyses. We employed the Cs+/$M^{+}$ configuration: positive primary ions with a Cs+ source and detection of positive secondary $M^{+}$ ions of mass $m$. The energy of the Cs+ primary beam is set to 10 keV. The raster size is 150$\times$150 $\mu$m2 with an analyzed zone restricted to a diameter of 33 $\mu$m and an electronic gate set to 70%. The analysis chamber is kept under high vacuum ($\approx$ 109 mbar). Secondary ions are detected in the positive mode by biasing the sample to +5000 V, leading to an interaction energy of the primary ions of 5 keV and an incidence angle of 46 with respect to the normal of the sample. We used a primary intensity of 10 nA. The depth of the resulting SIMS crater is measured with a Dektak8 step-meter. Knowing the total sputtering time for each oxide layer, we measured the sputtering rate in each of them. The values found for the Ba1-xSrxTiO3 layer, the La1.1Sr0.9NiO3 layer and the SrTiO3 substrate were of 0.25, 0.36, 0.24 nm/s respectively. Within the experimental resolution, no dependence of the sputtering rate was observed with the position in the compositionally graded layer. This excludes any large variations of the sputtering process, including the Cs+ primary ion implantation regime related to the changes in the Ba1-xSrxTiO3 composition.
@@
Due to the complexity of the sample layers, a particular care was given to the possible mass interferences. Indeed, the extraction energy was offset by +60 V in order to reduce the contribution of molecular ions as their intensities tend to decay faster in energy than the single ions [11]. Moreover, we verified that the mass spectra acquired in the compositionally graded layer actually match the natural relative abundance [13] of the selected elements (Figure 3). For example, the relative ionic intensities collected at integer masses from 46 to 50 corresponds to the natural abundance of 46Ti+, 47Ti+, 48Ti+, 49Ti+ and 50Ti+ which are 8.25%, 7.44%, 73.72%, 5.41% and 5.18% respectively [12]. This ensures that the contribution of single ions to the intensities collected at the analysed masses are significantly larger than the one of interfering molecular ions, if any. To minimize the complexity of the mass interferences, the analysed masses were chosen as small as possible and focused on the single ions 46Ti+, 48Ti+, 86Sr+, 88Sr+, 138Ba+ and 139La+.
@@ III.1 Matrix effect
Depth profiles for the selected masses were collected in the same conditions at evenly spaced positions along the chemical composition gradient; a stitching of optical images of the analyses craters is inserted in the inset of Figure 5. One of the greatest interests of the sample is the presence of the SrTiO3 substrate underneath compositionally graded layer. Its composition is close or identical to that of the Ba1-xSrxTiO3 top layer so it can be used as a constant reference along the gradient direction. Moreover, each reference is acquired in the same depth profile as the signal from the layer of interest. This makes the normalization procedure very robust and sound. As expected, no dependence of the ionic intensities from the substrate with the lateral position $X$ is observed in the substrate. The small variations are typical of irreproducibility of the instrument collecting factor as the lateral position is changed. The isotopic ratios of both titanium and strontium fit the natural abundance, thus confirming the absence of mass interferences at the analysed masses in the substrate too. In the following, those substrate 46Ti+ intensities will be used to normalize the intensities measured in the graded film. The normalized intensities of 88Sr+ and 138Ba+ measured in the Ba1-xSrxTiO3 layer are plotted against the $X$ position in Figure 5. The opposite trends of these too elements is in very good agreement with the chemical gradient as it was evidenced Ref. [10]. The intensities of Sr and Ba vary linearly with the Sr stoichiometric parameter $x$.
The situation for titanium is quite different. As Ba is substituted by Sr, the 48Ti+ signal level increases by about 25 % in the Ba1-xSrxTiO3 graded layer while it concentration remains constant. This evidences a well-characterised matrix effect that is a significant dependence of the titanium ionic signal with its chemical environment.

### v2

@@ I Introduction
This original growth technique offers an ideal ground for exploring the secondary ion mass spectrometry SIMS matrix effect. This effect designates variations of signal which are not related to the atomic concentration. The secondary ion intensity collected by the instrument indeed combines several factors including the concentration of the atom of interest in the bombarded material, its the sputtering yield and its ionization efficiency. The latter two depends on the chemical and morphological properties of the matrix. Therefore significant variations of the ionic intensity can occur from sample to sample independently to the atomic concentration which is further sought. Lots of efforts have been made to elucidate the effect of a matrix on the ionic yield of diluted traces. Although empirical trends have been identified very early [5], the vast variety of the material chemistry and of the analysis conditions have made the task tedious and little success to build a unified theory for all cases have been met so far [6]. A common approach to reach SIMS quantitative measurements consists of interpolating among chemically similar standards using semi-empirical models with limited validity domains [7]. This means to track ion yields in samples with similar compositions; it can be either series of alloys [8] or multi-implanted samples [9].
@@ II.2 SIMS conditions
The SIMS analyses were performed with a magnetic sector analyser (IMS7f-CAMECA). To avoid surface charging, the sample surface was coated with gold (50 nm) prior analyses. We employed the Cs+/$M^{+}$ configuration: positive primary ions with a Cs+ source and detection of positive secondary $M^{+}$ ions of mass $m$. The energy of the Cs+ primary beam is set to 10 keV. The raster size is 150$\times$150 $\mu$m2 with an analyzed zone restricted to a diameter of 33 $\mu$m and an electronic gate set to 70%. The analysis chamber is kept under high vacuum ($\approx$ 109 mbar). Secondary ions are detected in the positive mode by biasing the sample to +5 kV, leading to an interaction energy of the primary ions of 5 keV and an incidence angle of 46 with respect to the normal of the sample. We used a primary intensity of 10 nA. The depth of the resulting SIMS crater is measured with a Dektak8 step-meter. Knowing the total sputtering time for each oxide layer, we measured the sputtering rate in each of them. The values found for the Ba1-xSrxTiO3 layer, the La1.1Sr0.9NiO3 layer and the SrTiO3 substrate were of 0.25, 0.36, 0.24 nm/s respectively. Within the experimental resolution, no dependence of the sputtering rate was observed with the position in the compositionally graded layer. This excludes any large variations of the sputtering process, including the Cs+ primary ion implantation regime related to the changes in the Ba1-xSrxTiO3 composition.
@@
Due to the complexity of the sample layers, a particular care was given to the possible mass interferences. Indeed, the extraction energy was offset by -60 V in order to reduce the contribution of molecular ions as their intensities tend to decay faster in energy than the single ions [11]. Moreover, we verified that the mass spectra acquired in the compositionally graded layer actually match the natural relative abundance [13] of the selected elements (Figure 3). For example, the relative ionic intensities collected at integer masses from 46 to 50 corresponds to the natural abundance of 46Ti+, 47Ti+, 48Ti+, 49Ti+ and 50Ti+ which are 8.25%, 7.44%, 73.72%, 5.41% and 5.18% respectively [12]. This ensures that the contribution of single ions to the intensities collected at the analysed masses are significantly larger than the one of interfering molecular ions, if any. To minimize the complexity of the mass interferences, the analysed masses were chosen as small as possible and focused on the single ions 46Ti+, 48Ti+, 86Sr+, 88Sr+, 138Ba+ and 139La+.
@@ III.1 Matrix effect
Depth profiles for the selected masses were collected under the same conditions at evenly spaced positions along the chemical composition gradient; a stitching of optical images of the analyses craters is inserted in the inset of Figure 5. One of the greatest interests of the sample is the presence of the SrTiO3 substrate underneath compositionally graded layer. Its composition is close or identical to that of the Ba1-xSrxTiO3 top layer so it can be used as a constant reference along the gradient direction. Moreover, each reference is acquired in the same depth profile as the signal from the layer of interest. This makes the normalization procedure very robust and sound. As expected, no dependence of the ionic intensities from the substrate with the lateral position $X$ is observed in the substrate. The small variations are typical of irreproducibility of the instrument collecting factor as the lateral position is changed. The isotopic ratios of both titanium and strontium fit the natural abundance, thus confirming the absence of mass interferences at the analysed masses in the substrate too. In the following, those substrate 46Ti+ intensities will be used to normalize the intensities measured in the graded film. The normalized intensities of 88Sr+ and 138Ba+ measured in the Ba1-xSrxTiO3 layer are plotted against the $X$ position in Figure 5. The opposite trends of these two elements are in very good agreement with the chemical gradient as it was evidenced Ref. [10]. The intensities of Sr and Ba vary linearly with the Sr stoichiometric parameter $x$.
The situation for titanium is quite different. As Ba is substituted by Sr, the 48Ti+ signal level increases by about 25 % in the Ba1-xSrxTiO3 graded layer while it concentration remains constant. This evidences a well-characterised matrix effect that is a significant dependence of the titanium ionic signal on its chemical environment.